# Sharp thresholds for spanning regular subgraphs

### v2

@@ 6.8. Proof of Claim 34
where $o=o(x)\leq 8\sigma$ is chosen in such a way that (${x\choose o}$ achieves its maximum. Since $\left(\frac{e^{A_{1}+1}\sqrt{n}}{c}\right)^{c}\leq e^{A_{1}\sqrt{n}}$, we get that
$\left(\frac{e^{A_{1}+1}\sqrt{n}}{c}\right)^{c}\leq e^{A_{1}\sqrt{n}}\leq e^{(A_{1}/C)\ell}.$
Let us recall that $\ell>\ell_{0}$, and thus $x>\ell_{0}/2$. Here $\ell_{0}=\lfloor C\sqrt{n}\rfloor$, and $C\gg\max\{\frac{1}{\varepsilon^{\prime}},A_{1}/\delta\}$ can be chosen arbitrarily large.
@@ 2.2. If $c<\varepsilon^{\prime}x$ and $x\geq\varepsilon^{\prime}n$, then
since $\ell/2=x-3c/2-\sigma/2\geq x(1-5\varepsilon^{\prime}/2)$.
@@
$\leq n^{2}\sum_{\ell\geq\ell_{0}+1}\left(\frac{e^{\delta/2}+o(1)}{1+\varepsilon}\right)^{\ell}=o\left(\frac{1}{n}\right)$
@@
We then proceed in the same way as in the proof of Theorem 1.7. Here, $w=o(\ln n/r)$ and $\chi=\left\lfloor\frac{wn^{(2d-\Delta)/d}}{\ln n}\right\rfloor$. We replace diamonds with $(2r)$-subgraphs isomorphic to the union of two consecutive isomorphic copies of $F_{0}$, that is, $F_{0}^{*}\cong F[[2r]]$. Note that $F_{0}^{*}$ has $\lambda^{*}\geq 2\lambda+\Delta/2$ edges. In the same way as in Section 6, we define the family $\mathcal{F}_{n}(\overrightarrow{D})\subset\mathcal{F}_{n}$ for a $d$-tuple $\overrightarrow{D}$ of disjoint copies of $F_{0}^{*}$. We then fix $C>0$ large enough and sample an $m$-subset $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$ uniformly at random, where $m=\lfloor(1+\varepsilon)(e/n)^{2/d}\cdot N\rfloor$. We also let $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ be the set of all $F\setminus(D_{1}\cup\ldots\cup D_{\chi})$, $F\in\mathcal{F}_{n}(\overrightarrow{D})$. In the same way, for $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and an $m$-set $W\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$, we say that the pair $(F,W)$ is $\ell$-bad, if at least $n^{-1/2}$-fraction of $F^{\prime}\subset F\cup W$ that belong to $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ have $|F\cap F^{\prime}|>\ell$. We have that Claim 6.4 holds for these $\overrightarrow{D}$, $\mathcal{F}^{*}_{n}(\overrightarrow{D})$, $m$, and $\ell_{0}=\lfloor Cn^{(2d-\Delta)/d}\rfloor$, as well. Indeed, the proof of Claim 6.2 remains unchanged. Claim 6.4 follows then from 1) the fact that, for every $t\in\{0,1,\ldots,dn/2-\lambda^{*}\chi\}$,
@@
$\sum_{\ell\geq\ell_{0}+1}\Pi_{\ell}e^{-\frac{(dn/2)^{2}}{m}(1-o(1))}\left((1+o(1))\frac{N}{m}\right)^{\ell}=o\left(\frac{1}{n}\right).$
@@
$p(\ell,x,c)\leq\frac{{n\choose c}{x\choose c+\chi}\left(\frac{dx/2}{\chi}\right)^{\chi}e^{A_{1}c+A_{2}\sigma}\left(\max_{o\leq(d+4)\sigma}{x\choose o}\right)^{2}(n-x-2r\chi+c)!}{|\mathcal{F}_{n}(\overrightarrow{D})|}.$
@@
The key novel ingredient in the proof of Theorem 1.7 — the improvement of fragments, described in Section 6.2 — applies here as well. The main difference is that diamonds are replaced here with isomorphic copies of $F^{*}_{0}$. Let $\mu:=\left\lfloor\frac{10C\ln n}{rw}\right\rfloor$. We also consider all inclusion-maximal closed subgraphs of a typical $(F,\mathbf{W})$-fragment $H$ of size at most $\ell_{0}+\lambda^{*}\chi$ with at least $r\mu$ vertices, for $F\in\mathcal{F}_{n}(\overrightarrow{D})$. Here we cannot cut a closed subgraph in an arbitrary place since we have to keep pieces of size $r$ entirely. For each considered closed subgraph $P_{i}=(v^{i}_{1}v^{i}_{2}\ldots)$, we find the minimum $r^{\prime}$ such that $H[\{v^{i}_{r^{\prime}+1}\ldots v^{i}_{r^{\prime}+r}\}]\cong F_{0}$, the bijection preserving the order of vertices is an isomorphism, and $v^{i}_{r^{\prime}+1},\ldots,v^{i}_{r^{\prime}+r}\notin\partial_{v}(H)$. Then we remove from $H$ all the consecutive pieces $(v^{i}_{r^{\prime}+1}\ldots v^{i}_{r^{\prime}+r\mu})$, $(v^{i}_{r^{\prime}+r\mu+1}\ldots v^{i}_{r^{\prime}+2r\mu})$, $\ldots$, of size $r\mu$. We also have to make sure that all vertices of the last removed piece are not boundary. Finally, we glue the two remaining pieces of every $H$, and insert each removed piece between the two isomorphic copies of $F_{0}$ of the respective $D_{i}\cong F_{0}^{*}$.
@@
This is typically possible due to Claim 6.5, that holds in this setting as well. For a bipartite graph $B$ with typical degrees, we consider the set $\Sigma=\Sigma(B)$ of all pairs $(F\in\mathcal{F}_{n},W\in{{[n]\choose 2}\choose m})$ such that there exists $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D}(F))$ satisfying
@@
    $B[V^{\prime}\times U]$ has a matching that covers all vertices in $U$ that represent positions of first vertices of the removed pieces of the fragment $H=F\cap F^{\prime}$.
Due to symmetry and linearity of expectation, we get that there exists $B$ such that (asymptotically) almost all pairs $(F,W)$ belong to $\Sigma$. We fix such $B$. In order to complete analysis of smoothed fragments, we state the following analogue of Claim 6.7.
In the same way as in Section 6.3, for every pair $(F,W)\in\Sigma$, we define the distribution $\mathcal{Q}_{F,W}$ over the set of $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D}(F))$ such that $F^{\prime}\subset F\cup W$. We fix any positive $\delta<1/d$ and call the pair $(F,W)$ separating, if for every set $\mathcal{X}$ of disjoint closed subgraphs of $F$ on at least $n^{(2d-\Delta)/d}/\ln n$ vertices, the probability that the set of inclusion-maximal closed subgraphs of $\mathbf{F}^{\prime}\cap F$, where $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$, contains $\mathcal{X}$ is at most $\left\lfloor\frac{\ln^{2}n}{n^{\delta}}\right\rfloor$. For $W\in{{[n]\choose 2}\choose m}$, let $\Sigma_{W}$ be the set of all $F\in\mathcal{F}_{n}$ such that $(F,W)\in\Sigma$ is separating. For every $F\in\Sigma_{W}$, we pick $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$ independently. Let $H(F):=F\cap\mathbf{F}^{\prime}$ and $H^{\prime}(F)$ be the smoothed version of $H(F)$.
@@
$|\mathcal{S}(S)|\leq r^{c(S)}(16d)^{\ell(S)}(n-x(S)+c(S))!\max\left\{n^{c(S)(\Delta-d)/(3d)},{\ell_{0}+\lambda^{*}\chi\choose c(S)+\sigma(S)}\right\}.$
Moreover, the bound
$|\mathcal{S}(S)|\leq r^{c(S)}(16d)^{\ell(S)}(n-x(S)+c(S))!\frac{(\ell_{0}+\lambda^{*}\chi)^{c(S)+\sigma(S)}}{\min\left\{(\ln n)^{c(S)+\sigma(S)},n^{1/(2d)}\right\}}$
holds with probability $1-e^{-\Omega(n^{1/6})}$.
@@
Then, as in Section 6.5, we perform one fragmentation step and reduce the size of a fragment by the factor of $\Omega\left(n^{\min\{\frac{2\Delta-2d}{3d},\frac{2d-\Delta}{2d}\}-\delta}\right)$, for an arbitrarily small $\delta=\delta(d)>0$. Indeed, let $\mathcal{H}^{(1)}$ be the set of smoothed fragments, let $\mathbf{H}$ be a uniformly random element of $\mathcal{H}^{(1)}$, and let $\mathbf{W}^{(1)}\sim G(n,m^{\prime})$, $m^{\prime}=\lfloor\varepsilon n^{-2/d}\cdot N\rfloor$, be sampled independently.
@@
Whp there exists $H\subset\mathbf{H}\cup\mathbf{W}^{(1)}$ such that $H\in\mathcal{H}^{(1)}$ and
$|\mathbf{H}\cap H|\leq n^{\frac{(2d-\Delta)}{d}-\min\{\frac{2\Delta-2d}{3d},\frac{2d-\Delta}{2d}\}+\delta}.$
In order to prove this claim, we apply Lemma 4.1 with $f=\ell_{0}+\lambda^{*}\chi$ and $\mathcal{B}$ being the set of all graphs on $n$ vertices with more than $f^{(1)}:=n^{\frac{(2d-\Delta)}{d}-\min\{\frac{2\Delta-2d}{3d},\frac{2d-\Delta}{2d}\}+\delta}$ edges. It is then sufficient to show that, uniformly in $H\in\mathcal{H}^{(1)}$,
@@
$\sum_{\ell\geq f^{(1)}+1}\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\sum_{c,\ell,x}\left[\frac{{\ell_{0}+\lambda^{*}\chi\choose c}n^{(\Delta-d)c/(3d)}}{n^{(\Delta-d)c/d-\delta c}}e^{-C^{\prime}\ell}+\left(\frac{(\ell_{0}+\lambda^{*})n^{\delta}}{(c+\sigma)^{2}}\right)^{c+\sigma}\right]$
@@ 8. Remaining challenges
A notable special case of graphs $F$ covered by this conjecture is a hexagonal lattice. Although the hexagonal lattice is not a regular graph, it can be extended to a 3-regular graph by adding missing edges between the boundary vertices of degree 2. Since the boundary has $O(\sqrt{n})$ vertices, it does not contribute to the threshold asymptotically. This specific $F$ was mentioned by Riordan in [31], where he considered it as a special case to which his methods do not apply. It is worth noting that Fernandez de la Vega and Manoussakis [10] proved that $p_{c}(F)=O((\log n/n)^{2/3})$ and that Theorem 1.2, as well as [7, Corollary 1.5], implies (1) for such $F$.
It is also interesting to establish stronger versions of Theorems 21.7, and 1.8 for $d$-regular graphs $F$ that they cover. Although it is plausible to refine the asymptotics of $p_{c}(F)$ to some extent using similar methods, achieving results as precise as those for Hamilton cycles seems very challenging. Let us recall that, for an $n$-cycle $F$, $p_{c}(F)=\frac{\ln n+\ln\ln n+O(1)}{n}$ [20] and that the following hitting time result is known [1, 4]: in the random graph process $\mathbf{W}=\mathbf{W}_{m}$, $m=0,1,\ldots,N$, the hitting time for $\mathcal{Q}_{F}$ coincides with the hitting time for $\delta(\mathbf{W})\geq 2$ whp. It is clear that for denser regular graphs $F$ hitting time versions of Theorems 21.7, and 1.8 do not hold: for example, it is possible to show that, for any $k=(1-\Theta(1))n$, whp the hitting time for containing the second power of a Hamilton cycle is strictly greater than the hitting time for the property that every vertex belongs to the second power of a path of length at least $k$. On the other hand, our proof method suggests that the following question may have a positive answer: Is there a (small) constant $k\in\mathbb{N}$ such that $p_{c}(F)\leq p_{e}(F)+p_{c}(\mathcal{Q}_{k})$, where $F$ is the second power of a cycle on $[n]$ and $\mathcal{Q}_{k}$ denotes the property that every vertex belongs to the second power of a path of length $k$? In particular, is it true for $k=2$, i.e. when $\mathcal{Q}_{k}$ is the property that there are no isolated vertices?

### v3

@@ 6.8. Proof of Claim 39
where $o=o(x)\leq 8\sigma$ is chosen in such a way that (${x\choose o}$ achieves its maximum. Since $\left(\frac{e^{A_{1}+1}\sqrt{n}}{c}\right)^{c}\leq e^{e^{A_{1}}\sqrt{n}}$, we get that
$\left(\frac{e^{A_{1}+1}\sqrt{n}}{c}\right)^{c}\leq e^{e^{A_{1}}\sqrt{n}}\leq e^{(e^{A_{1}}/C)\ell}.$
Let us recall that $\ell>\ell_{0}$, and thus $x>\ell_{0}/2$. Here $\ell_{0}=\lfloor C\sqrt{n}\rfloor$, and $C\gg\max\{\frac{1}{\varepsilon^{\prime}},e^{A_{1}}/\delta\}$ can be chosen arbitrarily large.
@@ 2.2. If $c<\varepsilon^{\prime}x$ and $x\geq\varepsilon^{\prime}n$, then
since $\ell/2=x-3c/2-\sigma/2\geq x(1-2\varepsilon^{\prime})$.
@@
$\leq n^{2}\sum_{\ell\geq\ell_{0}+1}\left(\frac{e^{\delta/2}+o(1)}{1+\varepsilon}\right)^{\ell}=e^{-\Omega(\sqrt{n})},$
@@
We then proceed in the same way as in the proof of Theorem 1.7. Here, $w=o(\ln n/r)$ and $\chi=\left\lfloor\frac{wn^{(2d-\Delta)/d}}{\ln n}\right\rfloor$. We replace diamonds with $(2r)$-subgraphs isomorphic to a union of two consecutive isomorphic copies of $F_{0}$, that is, $F_{0}^{*}\cong F[[2r]]$. Note that $F_{0}^{*}$ has $\lambda^{*}\geq 2\lambda+\Delta/2$ edges. In the same way as in Section 6, we define the family $\mathcal{F}_{n}(\overrightarrow{D})\subset\mathcal{F}_{n}$, for a $d$-tuple $\overrightarrow{D}$ of disjoint copies of $F_{0}^{*}$. We then fix $C>0$ large enough and sample an $m$-subset $\mathbf{W}\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$ uniformly at random, where $m=\lfloor(1+\varepsilon)(e/n)^{2/d}\cdot N\rfloor$. We also let $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ be the set of all $F\setminus(D_{1}\cup\ldots\cup D_{\chi})$, $F\in\mathcal{F}_{n}(\overrightarrow{D})$. In the same way, for $F\in\mathcal{F}^{*}_{n}(\overrightarrow{D})$ and an $m$-set $W\subset{[n]\choose 2}\setminus(D_{1}\cup\ldots\cup D_{\chi})$, we say that the pair $(F,W)$ is $\ell$-bad, if at least $n^{-1/2}$-fraction of $F^{\prime}\subset F\cup W$ that belong to $\mathcal{F}^{*}_{n}(\overrightarrow{D})$ have $|F\cap F^{\prime}|>\ell$. We have that Claim 6.4 holds for these $\overrightarrow{D}$, $\mathcal{F}^{*}_{n}(\overrightarrow{D})$, $m$, and $\ell_{0}=\lfloor Cn^{(2d-\Delta)/d}\rfloor$, as well. Indeed, the proof of Claim 6.2 remains unchanged. Claim 6.4 follows then from 1) the fact that, for every $t\in\{0,1,\ldots,dn/2-\lambda^{*}\chi\}$,
@@
$\sum_{\ell\geq\ell_{0}+1}\Pi_{\ell}e^{-\frac{(dn/2)^{2}}{m}(1-o(1))}\left((1+o(1))\frac{N}{m}\right)^{\ell}=\exp(-\Omega(n^{2/d})).$
@@
$p(\ell,x,c)\leq\frac{{n\choose c}{x\choose c+\chi}\left(\frac{dx+2\chi}{\chi}\right)^{\chi}e^{A_{1}c+A_{2}\sigma}\left(\max_{o\leq(d+4)\sigma}{x\choose o}\right)^{2}(n-x-2r\chi+c)!}{|\mathcal{F}_{n}(\overrightarrow{D})|}.$
@@
In the same way as in Section 6.3, for every pair $(F\in\mathcal{F}_{n},W\in{{[n]\choose 2}\choose m})$, we define the distribution $\mathcal{Q}_{F,W}$ over the set of $F^{\prime}\in\mathcal{F}_{n}(\overrightarrow{D}(F))$ such that $F^{\prime}\subset F\cup W$. We fix any positive $\delta<1/d$ and call the pair $(F,W)$ separating, if for every set $\mathcal{X}$ of disjoint closed subgraphs of $F$ on at least $n^{(2d-\Delta)/d}/\ln n$ vertices, the probability that the set of inclusion-maximal closed subgraphs of $\mathbf{F}^{\prime}\cap F$, where $\mathbf{F}^{\prime}\sim\mathcal{Q}_{F,W}$, contains $\mathcal{X}$ is at most $\left\lfloor\frac{\ln^{2}n}{n^{\delta}}\right\rfloor$. We then generate $\mathbf{F}^{\prime}(F,W)\sim\mathcal{Q}_{F,W}$ for every pair $(F,W)$ and get fragments $F\cap\mathbf{F}^{\prime}(F,W)$.
The key novel ingredient in the proof of Theorem 1.7 — the improvement of fragments, described in Section 6.2 — applies here as well. The main difference is that diamonds are replaced here with isomorphic copies of $F^{*}_{0}$. Let $\mu:=\left\lfloor\frac{10C\ln n}{rw}\right\rfloor$. We also consider all inclusion-maximal closed subgraphs of a fragment $H=H(F,W)$ of size at most $\ell_{0}+\lambda^{*}\chi$ with at least $r\mu$ vertices, for $F\in\mathcal{F}_{n}(\overrightarrow{D})$. Here we cannot cut a closed subgraph in an arbitrary place since we have to keep pieces of size $r$ entirely. For each considered closed subgraph $P_{i}=(v^{i}_{1}v^{i}_{2}\ldots)$, we find the minimum $r^{\prime}$ such that $H[\{v^{i}_{r^{\prime}+1}\ldots v^{i}_{r^{\prime}+r}\}]\cong F_{0}$, the bijection preserving the order of vertices is an isomorphism, and $v^{i}_{r^{\prime}+1},\ldots,v^{i}_{r^{\prime}+r}\notin\partial_{v}(H)$. Then we remove from $H$ all the consecutive pieces $(v^{i}_{r^{\prime}+1}\ldots v^{i}_{r^{\prime}+r\mu})$, $(v^{i}_{r^{\prime}+r\mu+1}\ldots v^{i}_{r^{\prime}+2r\mu})$, $\ldots$, of size $r\mu$. We also have to make sure that all vertices of the last removed piece are not boundary. Finally, we glue the two remaining pieces of every $P_{i}$, and insert each removed piece between the two isomorphic copies of $F_{0}$ of the respective $D_{i}\cong F_{0}^{*}$.
@@
This is typically possible due to Claim 6.5, that holds in this setting as well. For a bipartite graph $B$ on $V\times U$ with typical degrees, we consider the (random) set $\Sigma=\Sigma(B)$ of all separating pairs $(F\in\mathcal{F}_{n},W\in{{[n]\choose 2}\choose m})$ such that $\mathbf{F}^{\prime}=\mathbf{F}^{\prime}(F,W)\sim\mathcal{Q}_{F,W}$ satisfies the following
@@
    $B[V^{\prime}\times U]$ has a matching that covers all vertices in $U$ that represent positions of first vertices of the removed pieces of the fragment $H=F\cap\mathbf{F}^{\prime}$.
Due to symmetry and linearity of expectation, we get that there exists $B$ such that whp the uniformly random vector $(\mathbf{F},\mathbf{W})$ belongs to $\Sigma$, and fix such $B$. In order to complete analysis of smoothed fragments, we state the following analogue of Claim 6.7. We fix $W$, let $H(F):=F\cap\mathbf{F}^{\prime}$, and let $H^{\prime}(F)$ be the smoothed version of $H(F)$, according to $B$.
@@
$|\mathcal{S}(S)|\leq r^{c}(16d)^{\ell}(n-x+c)!\max\left\{n^{c(\Delta-d)/(3d)}{\ell_{\chi}\choose\min\{\sigma,\lfloor\ell_{\chi}/2\rfloor\}},{\ell_{\chi}\choose\min\{c+\sigma,\lfloor\ell_{\chi}/2\rfloor\}}\right\}$
almost surely. Moreover, the bound
$|\mathcal{S}(S)|\leq r^{c}(16d)^{\ell}(n-x+c)!\frac{(2\ell_{\chi})^{c+\sigma}}{\min\left\{(\ln n)^{c+\sigma},n^{1/(2d)}\right\}}$
holds with probability $1-e^{-\Omega(n^{1/(2d)})}$.
@@
Then, as in Section 6.5, we perform one fragmentation step and reduce the size of a fragment by the factor of $\Omega\left(n^{\frac{\Delta-d}{2d}-\delta}\right)$, for any $\delta>0$. Indeed, let $\mathcal{H}^{(1)}$ be the set of smoothed fragments, let $\mathbf{H}$ be a uniformly random element of $\mathcal{H}^{(1)}$, and let $\mathbf{W}^{(1)}\sim G(n,m^{\prime})$, $m^{\prime}=\lfloor\varepsilon n^{-2/d}\cdot N\rfloor$, be sampled independently.
@@
Let $\delta>0$. Whp there exists $H\subset\mathbf{H}\cup\mathbf{W}^{(1)}$ such that $H\in\mathcal{H}^{(1)}$ and
$|\mathbf{H}\cap H|\leq n^{\frac{2d-\Delta}{d}-\frac{\Delta-d}{2d}+\delta}.$
In order to prove this claim, we apply Lemma 4.1 with $f=\ell_{0}+\lambda^{*}\chi$ and $\mathcal{B}$ being the set of all graphs on $n$ vertices with more than $f^{(1)}:=n^{\frac{(2d-\Delta)}{d}-\frac{\Delta-d}{2d}+\delta}$ edges. It is then sufficient to show that, uniformly in $H\in\mathcal{H}^{(1)}$,
@@
$\sum_{\ell\geq f^{(1)}+1}\Pi^{H}_{\ell}\left((1+o(1))\frac{N}{m^{\prime}}\right)^{\ell}$ $\leq\sum_{c,\ell,x}\left[\frac{{\ell_{0}+\lambda^{*}\chi\choose c}n^{(\Delta-d)c/(3d)}}{n^{(\Delta-d)c/d-\delta c}}e^{-C^{\prime}\ell}+\left(\frac{(\ell_{0}+\lambda^{*}\chi)^{2}n^{\delta/2}}{n^{(\Delta-d)/d}(c+\sigma)^{2}}\right)^{c+\sigma}\right]$
@@ 8. Remaining challenges
A notable special case of graphs $F$ covered by this conjecture is a hexagonal lattice. Although the hexagonal lattice is not a regular graph, it can be extended to a 3-regular graph by adding missing edges between the boundary vertices of degree 2. Since the boundary has $O(\sqrt{n})$ vertices, it does not contribute to the threshold asymptotically. This specific $F$ was mentioned by Riordan in [33], where he considered it as a special case to which his methods do not apply. It is worth noting that Fernandez de la Vega and Manoussakis [11] proved that $p_{c}(F)=O((\log n/n)^{2/3})$ and that Theorem 1.2, as well as [8, Corollary 1.5], implies (1) for such $F$.
It is also interesting to establish stronger versions of Theorems 21.7, and 1.8 for $d$-regular graphs $F$ that they cover. Although it is plausible to refine the asymptotics of $p_{c}(F)$ to some extent using similar methods, achieving results as precise as those for Hamilton cycles seems very challenging. Let us recall that, for an $n$-cycle $F$, $p_{c}(F)=\frac{\ln n+\ln\ln n+O(1)}{n}$ [22] and that the following hitting time result is known [1, 4]: in the random graph process $\mathbf{W}=\mathbf{W}_{m}$, $m=0,1,\ldots,N$, the hitting time for $\mathcal{Q}_{F}$ coincides with the hitting time for $\delta(\mathbf{W})\geq 2$ whp. It is clear that for denser regular graphs $F$ hitting time versions of Theorems 21.7, and 1.8 do not hold: for example, it is possible to show that, for any $k=(1-\Theta(1))n$, whp the hitting time for containing the square of a Hamilton cycle is strictly greater than the hitting time for the property that every vertex belongs to the square of a path of length at least $k$. On the other hand, our proof method suggests that the following question may have a positive answer: Is there a (small) constant $k\in\mathbb{N}$ such that $p_{c}(F)\leq p_{e}(F)+p_{c}(\mathcal{Q}_{k})$, where $F$ is the square of a cycle on $[n]$ and $\mathcal{Q}_{k}$ denotes the property that every vertex belongs to the square of a path of length $k$? In particular, is it true for $k=2$, i.e. when $\mathcal{Q}_{k}$ is the property that there are no isolated vertices?